OpenStudy (anonymous):

Which expression is equivalent to two-thirds divided by five-sixths? five-sixths times two-thirds two-thirds times five-sixths three-halves times six-fifths two-thirds times six-fifths

OpenStudy (lgbasallote):

hint: \[\huge \frac 34 \div \frac 23 \implies \frac 34 \times \frac 32\] does that help?

OpenStudy (anonymous):

no not really...

OpenStudy (shane_b):

That's an example...it's not your question. Just look at it to solve the problem you posted.

OpenStudy (anonymous):

but idk how to solve it...

OpenStudy (shane_b):

If you do the same thing that @lgbasallote posted with \[\Large \frac{2}{3} \div \frac{5}{6}\]what do you get? You just invert the second fraction and multiply...

OpenStudy (anonymous):

do i divide the top and bottom or just the top

OpenStudy (shane_b):

The question isn't asking you to solve the problem...it's asking you to convert it to multiplication. There's no division involved with just doing that. Just flip the second fraction and multiply....you'll see that what you come up with matches one of the answers.

OpenStudy (anonymous):

i am sooo lost right now

OpenStudy (shane_b):

Ok...I'll try to draw it. It's very simple

OpenStudy (anonymous):

ok

OpenStudy (shane_b):